# LOVA3: Learning to Visual Question Answering, Asking and Assessment

### v1

@@ Abstract
Question answering, asking, and assessment are three innate human traits crucial for understanding the world and acquiring knowledge. By enhancing these capabilities, humans can more effectively utilize data, leading to better comprehension and learning outcomes. However, current Multimodal Large Language Models (MLLMs) primarily focus on question answering, often neglecting the full potential of questioning and assessment skills. To this end, we introduce LOVA3, an innovative framework named “Learning tO Visual Question Answering, Asking and Assessment,” designed to equip MLLMs with these additional capabilities. Our approach involves the creation of two supplementary training tasks GenQA and EvalQA, aiming at fostering the skills of asking and assessing questions in the context of images. To develop the questioning ability, we compile a comprehensive set of multimodal foundational tasks. For assessment, we introduce a new benchmark called EvalQABench, comprising 64,000 training samples (split evenly between positive and negative samples) and 5,000 validation and testing samples. We posit that enhancing MLLMs with the capabilities to answer, ask, and assess questions will improve their multimodal comprehension and lead to better performance. We validate our hypothesis by training an MLLM using the LOVA3 framework and testing it on 10 multimodal benchmarks. The results demonstrate consistent performance improvements, thereby confirming the efficacy of our approach.
@@ 1 Introduction
asking a question is very valuable and even more important than answering a question. Indeed, humans also acquire knowledge from learning to ask questions since it encourages individuals to engage more deeply with information, thereby enhancing problem-solving skills [74, 75, 24, 62]. In addition to asking questions, humans also improve themselves through self-evaluation: humans try to identify the correctness of the answer and thus are involved in a deep understanding of our diverse world [92, 72]. These three learning mechanisms not only play a vital role in the human intelligence learning process but also empower humans with the corresponding abilities, including answering, asking, and assessing abilities. Among them, the answering ability is necessary to handle QA-like tasks; the asking ability allows AI models to interact with humans or other AI models for necessary information; and the evaluation ability assesses the solution candidates given by humans or other models with many applications, e.g., the filtering synthetic data.
@@
Although innate to humans, apart from answering ability, two other learning mechanisms of asking and assessment remain formidable challenges for contemporary multimodal large language models (MLLMs). Even though some advanced MLLMs [45, 90, 4, 101] have even achieved remarkable proficiency in handling multimodal questions pertaining to mathematics [5], science [45], and commonsense knowledge [19]. However, the focus of most MLLMs [45, 19, 4, 90] predominantly revolves around visual question answering (VQA) tasks. As a result, as shown in Fig. 1, current MLLMs, e.g., the representative LLaVA-1.5 [45], suffer from inferior performance on asking questions and self-assess question-answer pairs (QA), which underscores their efficacy as problem-solvers and prohibits profound multimodal understanding.
To address these challenges, we introduce two essential tasks: GenQA and EvalQA, aiming at bolstering the intelligence and robustness of MLLMs. GenQA focuses on enabling the model to generate diverse question-answer (QA) pairs for images, thus equipping the MLLM with the capability to ask questions. We believe that if an MLLM can successfully generate QA pairs for challenging tasks, it indicates a higher level of problem-solving ability. Specifically, we define the GenQA task to include not only generic VQA (e.g., VQAv2 [29] and GQA [33]) but also Multi-Choice VQA (MC VQA), and Multi-Turn VQA (MT) to increase the variety of data formats. Additionally, we incorporate two challenging multimodal grounding tasks into the training process: Referring Expression Comprehension (REC) and Referring Expression Generation (REG). Learning to generate the data of these grounding tasks forces the MLLM to extract fine-grained visual cues from images, such as explicit object localization and compositional relationships. This, in turn, enhances the multimodal reasoning ability of MLLMs. During training, we gather the relevant datasets for these tasks and transform them into a generative format using our proposed instruction template. EvalQA, on the other hand, involves tasking the MLLM to predict the correctness of a given visual-question-answer triplet. Recognizing the absence of datasets specifically designed to assess VQA quality, we have developed a new benchmark called EvalQABench for evaluating VQA data. Rather than asking humans to label such a dataset, we propose a new automatic pipeline for data construction. This benchmark comprises training, validation, and test sets, with each VQA pair accompanied by a “Yes” or “No” label indicating correctness, along with a one-sentence explanation as the feedback. For instance, “Yes, the oranges are not in a bag”.
@@
    We build a new benchmark EvalQABench for the VQA evaluation as the first effort to advance the development of future research.
@@
    The experimental results demonstrate that training with LOVA3 consistently improves performance across several multimodal benchmarks, including MME, VizWiz, and MMBench.
@@ 2.1 Multimodal Large Language Models
Large Language Models (LLMs)  [18, 104, 70, 7, 87, 17, 88] such as GPT-4 [66] demonstrate their exceptional capacity to handle a wide range of complex tasks to play an important role in assisting humans in daily life. Equipped with these LLMs, a surge of multimodal modes [41, 46, 19, 4, 12, 90, 108, 11, 10, 55, 91, 32, 67, 63, 94, 25, 50, 31, 38, 9, 106, 51, 16, 42, 27, 73, 26, 71] are proposed to integrate the visual information with the pre-trained LLM decoder for diverse multimodal reasoning tasks such as image captioning [14, 1, 97] and visual question answering [29, 59, 33, 30]. LLaVA [46, 45] is a pioneering approach that proposes to project the CLIP [69] visual features to the language embedding space and showcase promising performance on various multimodal benchmarks. Unlike the architecture of LLaVA, InstructBLIP [19] employs an instruction-aware feature extractor and obtains advanced performance on various tasks. Besides focusing on traditional vision-language tasks, Shikra [12], Kosmos-2 [67], PVIT [10], Ferret [96] pay attention to the image-region based multimodal tasks (i.e., Referring Expression Comprehension). By adopting a large-scale image-text corpus for instruction tuning, Qwen-VL [4], CogVLM [90], AnyMAL [63] and Chameleon [85] achieve exceptional performance on various multimodal tasks. However, these MLLMs primarily concentrate on training the model to answer questions as effectively as possible, neglecting the significance of enabling the model to act as a questioner and a competent evaluator within the training paradigm.
@@ 2.2 Visual Question Answering and Generation
Nine years ago, visual question answer [3] was defined and became an essential task for evaluating multimodal systems. A surge of VQA-related benchmarks  [3, 29, 82, 61, 60, 30, 76, 59, 54, 33, 43] are emerged to advance the development of this research area, including generic VQA benchmarks [3, 29, 33], text-based VQA [82, 61, 60, 84], knowledge-augmented VQA [59, 76, 54, 15], and benchmark [30] aimed at assisting blind people. Simutaneously, various studies [56, 80, 52, 23, 107, 6, 2, 35] are proposed with diverse model architecture such as LSTMs[80], attention network[35, 2] or convolution network[52, 107].
Besides the VQA task, the Visual Question Generation (VQG) task was first formulated in  [65], which contributes a VQG dataset with each image annotated with multi-questions and benchmarking on generative models and retrieval models.  [103] first employs an RNN-based encoder-decoder framework alongside model-generated captions to generate questions. After that a list of works [64, 65, 20, 36, 77, 93, 89] are proposed for promoting this research area. Two interesting studies [44, 78] pose that treating VQG as a complementary task can enhance the robustness of visual question answering. This finding reaffirms our motivation that training a model to generate diverse questions contributes to a deeper understanding of visual information, thereby improving its problem-solving capabilities.
As for assessing the quality of VQA triplets, to the best of our knowledge, there are no efforts to do that. Therefore, we propose a new task EvalQA for the model training tailored with a new benchmark EvalQABench to advance this research area.
@@ 2.3 Multimodal Benchmarks
Traditional multimodal benchmarks focus on answering ability, such as visual question answering [29], image captioning [14, 68, 1], as well as other benchmarks for specialized scenarios such as scene text understanding [82, 81], commonsense reasoning [100], outside knowledge [59, 76]. The recent development of MLLM posts a strong need for modernized multimodal benchmarks [21, 48, 40, 98, 28, 99, 53, 22, 13, 102, 47, 105, 83, 86] such as MME [21], MMBench [48], SEED-Bench [40] which involve comprehensively evaluating current MLLMs on various multimodal abilities.
Unlike existing multimodal benchmarks focusing primarily on evaluating the model’s ability to answer, we introduce EvalQABench, a benchmark designed to evaluate the quality of VQA pairs, each with a binary “Yes/No” annotation. Furthermore, recognizing the lack of emphasis on providing feedback for incorrect answers in current benchmarks, we develop an LLM-based pipeline. This pipeline can automatically generate feedback, paving the way for enhanced automated data processing in the future.
@@ 3.1 Data Collection for GenQA
If one MLLM is able to successfully generate high-quality question-answer pairs based on visual input, it indicates a stronger problem-solving ability and deep visual understanding. To enable the MLLM to ask questions, it is natural for us to gather existing annotated datasets as the training corpus and then train the model to predict both questions and answers. We carefully define five main multimodal data types as listed in Tab. 1. For each data type, we gather widely used human-annotated datasets or high-quality instruction tuning datasets generated by GPT-4. We select Generic VQA tasks to generate fundamental questions, e.g., object count and object action. We incorporate Multi-choice VQA (MC VQA) and Multi-turn VQA (MT) to increase the diversity of data formats. Additionally, we include two multimodal grounding tasks: Referring Expression Comprehension (REC) and Referring Expression Generation (REG). Generating REC and REG data requires a deeper understanding of image content, enabling the model to fully comprehend visual cues. Both tasks increase the difficulty of GenQA, which helps MLLM acquire a higher level of multimodal understanding. In total, we gather 842K data for training questioning ability.
@@ 3.2 Data Creation for EvalQA
MLLM-based Negative Answer Generation. The main challenge of EvalQABench lies in constructing negative answers. When dealing with large-scale ground-truth VQA pairs, how can we automatically produce the negative answer? One viable solution is to leverage a multimodal model for this purpose. Recognizing that Fuyu-8B [5] is an open-source free MLLM that stands out with the exceptional ability to process high-resolution images and perform robust well on many complex tasks, rivaling the capabilities of GPT-4V. We utilize it to generate negative answers with the following prompt:
@@
LLM-based Feedback Generation. With the candidate’s negative answer, we then focus on generating error feedback. We consider the feedback describing the reason for incorrectness will help the MLLM obtain a deeper understanding. We thus utilize the SOTA LLM Llama 2 [88] to generate the feedback by reasoning the ground truth question-answer pairs with the following prompt:
@@ 4.1 Datasets and Settings
Training Datasets. For the fair comparison, we utilize the 665K instruction-following dataset introduced in LLaVA1.5, combined with the 842K GenQA data as outlined in Tab. 1, and an additional 64K data comprising one-positive-one-negative pairs as described in Section 3.2, totaling our training datasets. It is important to note that the datasets and annotations used in both VQA and GenQA are the same. There are no additional datasets involved, thus avoiding unfair comparisons caused by the introduction of new instruction data. For EvalQA, we adopt VQAv2 to build the training set, which is already included in the original 665K instruction dataset.
@@
Implementation Details. To ensure a fair comparison, we train LOVA3 model without tuning any hyperparameters of LLaVA-1.5 [45] from its original supervised finetuing stage. The model is trained for one epoch across three tasks: VQA, GenQA, and EvalQA. Specifically, we employ the AdamW [49] optimizer with a learning rate of $2\times 10^{-5}$ and a total batch size of 128. The training process takes 24.5 hours on an 8 Nvidia A100 (40G) GPU setup.
@@
Method Train Paradigm LLM
VQAv2
test-dev
GQA
test
VizWiz
test-dev
ScienceQA
img
POPE
avg
@@
LOVA3(ours) VQA, GenQA, EvalQA Vicuna-7B 80.3${}^{*}_{+1.8}$ 63.3${}^{*}_{+1.3}$ 53.6+3.6 68.0+1.2 87.4+1.5
@@
LOVA3(ours) VQA, GenQA, EvalQA Vicuna-7B 1552.7+42.0 67.1+0.9 66.8+2.5 60.5+2.2 68.3+4.3
@@
LOVA3(ours) VQA, GenQA, EvalQA Vicuna-7B 41.5+4.5 23.6+2.6 23.9+6.3 24.6+4.2 30.3+5.4 35.2+4.0
@@ 4.2 Main Results
Generic tasks. As shown in Tab. 3, LOVA3 outperforms LLaVA1.5 across all five datasets and obtains 3.6% improvement on VizWiz dataset, 1.3% improvement on GQA, 1.8% improvement on VQAv2 (1,932 samples are correctly predicted), and 1.2% improvement on ScienceQA. As for the object hallucination benchmarks, our model attains 87.4% accuracy at an average of its three subsets. Remarkably, these enhancements in VQAv2 and GQA performance are achieved without any extra datasets, underscoring the significant impact of integrating GenQA and EvalQA into our training to promote performance improvements on these generic VQA tasks.
MME, SEED-Bench, MMBench, LLaVA-Bench. In Tab. 4, we evaluate four prevalent multimodal benchmarks, where our LOVA3 surpass LLaVA1.5 with 42.0% on MME benchmark, 0.9% increase in accuracy on SEED-Bench, 2.5% on MMBench (En), 2.2% MMBench (Cn) and 4.3% on LLaVA-Bench. Such results showcase enhanced multimodal reasoning capabilities for complex tasks compared to vanilla LLaVA1.5, which is solely trained with VQA tasks.
MM-Vet. In Tab. 5, we compare LOVA3 with other approaches on MM-Vet, which is a challenging benchmark including numerous complex VQA samples that demand integration of several multimodal capabilities for answering. As illustrated in Tab. 5, the results show that our LOVA3 outperforms LLaVA-1.5 by 4.0% at an average. Such improvement demonstrates the effectiveness of LOVA3 in solving these challenging multimodal questions.

### v2

@@ Abstract
Question answering, asking, and assessment are three innate human traits crucial for understanding the world and acquiring knowledge. By enhancing these capabilities, humans can more effectively utilize data, leading to better comprehension and learning outcomes. Current Multimodal Large Language Models (MLLMs) primarily focus on question answering, often neglecting the full potential of questioning and assessment skills. Inspired by the human learning mechanism, we introduce LOVA3, an innovative framework named “Learning tO Visual question Answering, Asking and Assessment,” designed to equip MLLMs with these additional capabilities. Our approach involves the creation of two supplementary training tasks GenQA and EvalQA, aiming at fostering the skills of asking and assessing questions in the context of images. To develop the questioning ability, we compile a comprehensive set of multimodal foundational tasks. For assessment, we introduce a new benchmark called EvalQABench, comprising 64,000 training samples (split evenly between positive and negative samples) and 5,000 validation and testing samples. We posit that enhancing MLLMs with the capabilities to answer, ask, and assess questions will enhance their multimodal comprehension, ultimately improving overall performance. To validate this hypothesis, we train MLLMs using the LOVA3 framework and evaluate them on a range of multimodal datasets and benchmarks. Our results demonstrate consistent performance gains, underscoring the critical role of these additional tasks in fostering comprehensive intelligence in MLLMs. The code is available at https://github.com/showlab/LOVA3.
@@ 1 Introduction
asking a question is very valuable and even more important than answering a question. Indeed, humans also acquire knowledge from learning to ask questions since it encourages individuals to engage more deeply with information, thereby enhancing problem-solving skills [69, 70, 21, 57]. In addition to asking questions, humans also improve themselves through self-evaluation: humans try to identify the correctness of the answer and thus are involved in a deep understanding of our diverse world [86, 67].
@@
Although innate to humans, apart from answering ability, two other learning mechanisms of asking and assessment remain underexplored for contemporary multimodal large language models (MLLMs). Current MLLMs [42, 84, 3, 95] are excelled in addressing diverse domains of multimodal questions such as mathematics [4], science [42], and commonsense knowledge [17]. However, they predominantly revolve around visual question answering (VQA). As a result, as shown in Fig. 1, current MLLMs, e.g., the representative LLaVA-1.5 [42], suffer from inferior performance on asking questions and self-assess question-answer pairs (QA), which underscores their efficacy as problem-solvers and prohibits holistic multimodal understanding.
To advance the comprehensive intelligence of MLLMs, we introduce two essential tasks: GenQA and EvalQA, aiming at bolstering the intelligence and robustness of MLLMs. GenQA focuses on enabling the model to generate diverse question-answer (QA) pairs from the single input image, thus equipping the MLLM with the capability to ask questions. We believe that if an MLLM can successfully generate QA pairs for challenging tasks, it indicates a higher level of problem-solving ability [40, 73]. Specifically, we define the GenQA task to include not only generic VQA (e.g., VQAv2 [26] and GQA [30]) but also Multi-Choice VQA (MC VQA), and Multi-Turn VQA (MT) to increase the variety of data formats. Additionally, we incorporate two challenging multimodal grounding tasks into the training process: Referring Expression Comprehension (REC) and Referring Expression Generation (REG). Learning to generate the data of these grounding tasks forces the MLLM to extract fine-grained visual cues from images, such as explicit object localization and compositional relationships. This, in turn, enhances the multimodal reasoning ability of MLLMs. During training, we gather the relevant datasets for these tasks and transform them into a generative format using our proposed instruction template. EvalQA, on the other hand, involves tasking the MLLM to predict the correctness of a given visual-question-answer triplet. Recognizing the absence of datasets specifically designed to assess VQA correctness, we have developed a new benchmark called EvalQABench for evaluating VQA data. Rather than asking humans to label such a dataset, we propose a new pipeline for data construction. This benchmark comprises training, validation, and test sets, with each VQA pair accompanied by a “Yes” or “No” label indicating correctness, along with a one-sentence explanation as the feedback. For instance, “Yes, the oranges are not in a bag”.
@@
    We build a new benchmark EvalQABench for the VQA evaluation as the first effort to advance the VQA data assessment of future research.
@@
    The experimental results demonstrate that training with LOVA3 consistently improves performance across several multimodal benchmarks, including VQAv2, GQA, MME, VizWiz, MMBench, and MM-Vet, etc.
@@ 2.1 Multimodal Large Language Models
Large Language Models (LLMs)  [16, 98, 65, 5, 81, 15, 82] such as GPT-4 [61] demonstrate their exceptional capacity to handle a wide range of complex tasks to play an important role in assisting humans in daily life. Equipped with these LLMs, a surge of multimodal modes [37, 43, 17, 3, 10, 84, 101, 9, 8, 51, 85, 29, 62, 58, 88, 22, 47, 28, 34, 7, 100, 48, 14, 38, 24, 68, 23, 66] are proposed to integrate the visual information with the pre-trained LLM decoder for diverse multimodal reasoning tasks such as image captioning [12, 1, 91] and visual question answering [26, 54, 30, 27]. LLaVA [43, 42] is a pioneering approach that collects 665K instruction tuning data from present vision-language (VL) datasets for supervised finetuning (SFT) and achieves promising results on various datasets in a lower cost of training requirement. Another SOTA model InstructBLIP [17] also proposes gathering datasets to construct their instruction tuning dataset. It adopts the VQG task but is limited in generic data type. Different from InstructBLIP, we propose the GenQA task for jointly generated questions and answers on 5 primary VL tasks not restricted to generic data type. Besides focusing on traditional vision-language tasks, Shikra [10], Kosmos-2 [62], PVIT [8], Ferret [90] pay attention to the image-region based multimodal tasks (i.e., Referring Expression Comprehension) and demonstrate the performance improvement with these hard tasks. By adopting a large-scale image-text corpus for instruction tuning, Qwen-VL [3], CogVLM [84], AnyMAL [58] and Chameleon [79] achieve exceptional performance on various multimodal tasks. However, these MLLMs primarily concentrate on training the model to answer questions as effectively as possible, neglecting the significance of enabling the model to act as a questioner and a competent evaluator within the training paradigm.
@@ 2.2 Visual Question Answering and Generation
Nine years ago, visual question answer [2] was defined and became an essential task for evaluating multimodal systems. A surge of VQA-related benchmarks  [2, 26, 76, 56, 55, 27, 71, 54, 50, 30, 39] are emerged to advance the development of this research area, including generic VQA benchmarks [2, 26, 30], text-based VQA [76, 56, 55, 78], knowledge-augmented VQA [54, 71, 50, 13], and goal-oriented VQA [27] aimed at assisting blind people.
Besides the VQA task, the Visual Question Generation (VQG) task was first formulated in  [60], which contributes a VQG dataset with each image annotated with multi-questions and benchmarking on generative models and retrieval models.  [97] first employs an RNN-based encoder-decoder framework alongside model-generated captions to generate questions. After that a list of works [59, 60, 18, 32, 72, 87, 83] are proposed for promoting this research area. Two interesting studies [40, 73] pose that treating VQG as a complementary task can enhance the robustness of visual question answering. This finding reaffirms our motivation that training a model to generate diverse questions contributes to a deeper understanding of visual information, thereby improving its problem-solving capabilities. Unlike the traditional Visual Question Generation (VQG) task, which focuses primarily on the generic VQA domain, GenQA is designed to generate diverse VQA data including MC VQA, MT, REC, and REG. Additionally, GenQA generates both questions and answers simultaneously, whereas traditional VQG focuses solely on question generation.
@@ 2.3 Multimodal Benchmarks
Traditional multimodal benchmarks focus on answering ability, such as visual question answering [26], image captioning [12, 63, 1], as well as other benchmarks for specialized scenarios such as scene text understanding [76, 75], commonsense reasoning [94], outside knowledge [54, 71]. The recent development of MLLM posts a strong need for modernized multimodal benchmarks [19, 45, 36, 92, 25, 93, 49, 20, 11, 96, 44, 99, 77, 80] such as MME [19], MMBench [45], SEED-Bench [36] which involve comprehensively evaluating current MLLMs on various multimodal abilities. Unlike existing multimodal benchmarks focusing primarily on evaluating the model’s ability to answer, we introduce EvalQABench, a benchmark designed to evaluate the correctness of VQA pairs, each with a binary “Yes/No” annotation. Furthermore, recognizing the lack of emphasis on providing feedback for incorrect answers in current benchmarks, we develop an LLM-based pipeline. This pipeline can automatically generate feedback, paving the way for enhanced automated data processing in the future.
@@ 3.1 Data Collection for GenQA
If one MLLM is able to successfully generate high-quality question-answer pairs based on visual input, it indicates a stronger problem-solving ability and deep visual understanding [40, 73]. To enable the MLLM to ask questions, it is natural for us to gather existing annotated datasets as the training corpus and then train the model to predict both questions and answers. We carefully define five main multimodal data types as listed in Tab. 1. For each data type, we gather widely used human-annotated datasets or high-quality instruction tuning datasets generated by GPT-4. We select Generic VQA tasks to generate fundamental questions, e.g., object count and object action. We incorporate Multi-choice VQA (MC VQA) and Multi-turn VQA (MT) to increase the diversity of data formats. Additionally, we include two multimodal grounding tasks: Referring Expression Comprehension (REC) and Referring Expression Generation (REG). Generating REC and REG data requires a deeper understanding of image content, enabling the model to fully comprehend visual cues. Both tasks increase the difficulty of GenQA, which helps MLLM acquire a higher level of multimodal understanding. In total, we gather 842K data for training questioning ability.
@@ 3.2 Data Creation for EvalQA
MLLM-based Negative Answer Generation. The main challenge of EvalQABench lies in constructing negative answers. When dealing with large-scale ground-truth VQA pairs, how can we automatically produce the negative answer? One viable solution is to leverage a multimodal model for this purpose. Recognizing that Fuyu-8B [4] is an open-source free MLLM that stands out with the exceptional ability to process high-resolution images and perform robust well on many complex tasks. We utilize it to generate negative answers with the following prompt:
@@
LLM-based Feedback Generation. With the candidate’s negative answer, we then focus on generating error feedback. We consider the feedback describing the reason for incorrectness will help the MLLM obtain a deeper understanding. We thus utilize the LLM Llama 2 [82] to generate the feedback by reasoning the ground truth question-answer pairs with the following prompt:
@@ 4.1 Datasets and Settings
Training Datasets. For the fair comparison, we utilize the 665K instruction-following dataset introduced in LLaVA1.5, combined with the 842K GenQA data as outlined in Tab. 1, and an additional 64K data comprising one-positive-one-negative pairs as described in Section 3.2, totaling our training datasets with 1.5M samples. It is important to note that the datasets and annotations used in both VQA and GenQA are the same. There are no additional datasets involved, thus avoiding unfair comparisons caused by the introduction of new instruction data. For EvalQA, we adopt VQAv2 to build the training set, which is already included in the original 665K instruction dataset.
@@
Implementation Details. To ensure a fair comparison, we train the LOVA3-7B model without tuning any hyperparameters of LLaVA-1.5 [42] from its original supervised finetuing stage. The model is trained for one epoch across three tasks: VQA, GenQA, and EvalQA. Specifically, we employ the AdamW [46] optimizer with a learning rate of $2\times 10^{-5}$ and a total batch size of 128. The training process takes 24.5 hours on an 8 Nvidia A100 (40G) GPU setup. Moreover, we also replace the LLM from Vicuna-7B[15] to Phi-1.5B[41] to evaluate smaller LLMs. We train LLaVA-Phi-1.5 and LOVA3-1.5B by using the same training recipe. The only difference of training with Phi-1.5 is that we increase the learning rate from $2\times 10^{-5}$ to $4\times 10^{-5}$ to ensure the higher performance. The model LLaVA-Phi-1.5 is trained with the original 665K VQA instruction data as the baseline. The model LOVA3-1.5B is trained with our proposed 1.5M mixture data including VQA, GenQA, EvalQA data.
@@
Method Train Paradigm LLM VQAv2 GQA VizWiz ScienceQA POPE
test-dev test test-dev img avg
LLaVA-Phi-1.5[42] VQA Phi-1.5B 73.2 56.1 33.8 57.3 87.6
[…]
@@
LOVA3-7B(ours) VQA, GenQA, EvalQA Vicuna-7B 80.3${}^{*}_{+1.8}$ 63.3${}^{*}_{+1.3}$ 53.6+3.6 68.0+1.2 87.4+1.5
@@
LOVA3-7B (ours) VQA, GenQA, EvalQA Vicuna-7B 1552.7+42.0 67.1+0.9 66.8+2.5 60.5+2.2 68.3+4.3
@@
LOVA3-7B (ours) VQA, GenQA, EvalQA Vicuna-7B 41.5+4.5 23.6+2.6 23.9+6.3 24.6+4.2 30.3+5.4 35.2+4.0
@@ 4.2 Main Results
Generic tasks. As shown in Tab. 3, LOVA3-7B outperforms LLaVA1.5 across all five datasets and obtains 3.6% improvement on VizWiz dataset, 1.3% improvement on GQA, 1.8% improvement on VQAv2 (1,932 samples are correctly predicted), and 1.2% improvement on ScienceQA. As for the object hallucination benchmarks, our model attains 87.4% accuracy at an average of its three subsets. Remarkably, these enhancements in VQAv2 and GQA performance are achieved without any extra datasets, underscoring the significant impact of integrating GenQA and EvalQA into our training to promote performance improvements on these generic VQA tasks. Based on the results from smaller LLMs, our LOVA3-1.5B outperforms the baseline LLaVA-Phi-1.5 on VQAv2, GQA, VizWiz, and ScienceQA by 2.6%, 2.5%, 3.4%, and 0.5%, respectively. This demonstrates a consistent improvement when training with our LOVA3 framework across varying LLM sizes. Additionally, a comparison of improvements for both 7B and Phi-15B on the VizWiz dataset highlights the advantage of our LOVA3 framework for this VQA task.
MME, SEED-Bench, MMBench, LLaVA-Bench. In Tab. 4, we evaluate four prevalent multimodal benchmarks, where our LOVA3-7B surpasses LLaVA1.5 with 42.0% on MME benchmark, 0.9% increase in accuracy on SEED-Bench, 2.5% on MMBench (En), 2.2% MMBench (Cn) and 4.3% on LLaVA-Bench. Such results showcase enhanced multimodal reasoning capabilities for complex tasks compared to vanilla LLaVA1.5, which is solely trained with VQA tasks. By investigating the results produced by LOVA3-1.5B on these multimodal benchmarks, one can see greater improvements in the smaller models. Notably, LOVA3-1.5B achieves an impressive 98.2% improvement on the MME benchmark. The lower results of LLaVA-Phi-1.5 and LOVA3-1.5B on MMBench (Cn) may be attributed to a lack of Chinese training data in the Phi-1.5 training process.
MM-Vet. In Tab. 5, we compare LOVA3-7B with other approaches on MM-Vet, which is a challenging benchmark including numerous complex VQA samples that demand integration of several multimodal capabilities for answering. As illustrated in Tab. 5, the results show that our LOVA3-7B outperforms LLaVA-1.5 by 4.0% at an average. Such improvement demonstrates the effectiveness of LOVA3-7B in solving these challenging multimodal questions. Based on the results on LOVA3-1.5B and LLaVA-Phi-1.5, one can see a greater improvement than LOVA3-7B.